# Permissive Information-Flow Analysis for Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) are rapidly becoming commodity components of larger software systems. This poses natural security and privacy problems: poisoned data retrieved from one component can change the model’s behavior and compromise the entire system, including coercing the model to spread confidential data to untrusted components. One promising approach is to tackle this problem at the system level via dynamic information flow (aka taint) tracking. Unfortunately, this approach of propagating the most restrictive input label to the output is too conservative for applications where LLMs operate on inputs retrieved from diverse sources.
In this paper, we propose a novel, more permissive approach to propagate information flow labels through LLM queries. The key idea behind our approach is to propagate only the labels of the samples that were influential in generating the model output and to eliminate the labels of unnecessary inputs. We implement and investigate the effectiveness of two variations of this approach, based on (i) prompt-based retrieval augmentation, and (ii) a $k$-nearest-neighbors language model. We compare these with a baseline that uses introspection to predict the output label. Our experimental results in an LLM agent setting show that the permissive label propagator improves over the baseline in more than 85% of the cases, which underscores the practicality of our approach.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 [32], Llama [41, 11], Mistral [18], and PaLM [1] are rapidly becoming commodity components of larger software systems. The inputs to these LLMs often consist of data retrieved from a variety of sources, including websites, productivity software, or tools [6], and their output is usually passed on to other software components for further processing [49].
This poses natural security and privacy problems: low integrity inputs (e.g., poisoned data) can change the model’s behavior in unexpected ways and potentially affect the entire system [15]. Similarly, high confidentiality inputs (e.g., confidential documents) can be inadvertently leaked to an untrusted downstream component [28].
One possible approach to address this problem is to rely on the LLM itself for mitigation, for example via introspection of the retrieved inputs or guardrails given in the meta-prompt. However, such defenses can be circumvented with more advanced attacks [54, 52, 23, 38], leading to the undesirable cat-and-mouse game that is common in system security. A promising alternative is to tackle this problem at the system level via dynamic information-flow (aka taint) tracking. Information-flow tracking is a standard technique for enforcing integrity and confidentiality properties [7, 5, 35], and has been used successfully in many applications, including detecting cross-site-scripting vulnerabilities [43], privacy leaks in mobile applications [12] and recently to LLM-based systems [48, 47, 2].
In information-flow tracking, each piece of data is augmented with labels describing its integrity or confidentiality. Labels are usually propagated conservatively: the output of an operation on data is labeled as the most restrictive (i.e., most confidential or least trusted) label of its inputs. For example, the output of a function that takes two arguments, one trusted and one untrusted, would be labeled as untrusted. A challenge of applying such label propagation mechanisms to LLMs is that the output label would be the upper bound of all inputs (i.e., the context) used for inference. With LLMs having the ability to retrieve documents from different sources, this can quickly become unnecessarily restrictive, a phenomenon known in the literature as label creep [35].
@@
To ground our work in a relevant LLM system, we implement and evaluate two different realizations of our label propagator: (i) a prompt-based system [22] where the retrieved documents are provided within the prompt to an autoregressive LLM, and (ii) a $k$NN-LM architecture [20], where the output distribution is computed as a mixture of the distribution of the model and the retrieved documents.
@@ 2 Problem Setting and Goal
This formulation is sufficiently general to represent many real-world applications of LLMs. For example, in retrieval-augmented generation (RAG) [22, 16], the context contains documents retrieved from the knowledge-base. Alternatively, if an LLM uses plugins to retrieve external data (e.g., web search, email retrieval, calendar query, etc.) [36], the data returned by the plugins is part of the context.
@@ 3.1 $\lambda$-similar Labels
Our goal is to find labels $L^{\prime}\sqsubseteq L$ such that the output of the language model changes only negligibly when substituting $C$ with $C_{|L^{\prime}}$. We capture "negligible change" by introducing a hyperparameter $\lambda$ and require that the utility of the model’s output under the full context drops by at most $\lambda$ when restricting to the subcontext. Formally:
@@
Definition 1 leaves the choice of the utility function $U$ and the language model $p_{\text{LM}}(y|x,C)$ unspecified, as different applications require custom choices of these functions. In this paper, we focus on language modeling where perplexity is a common way to measure utility [19]. Hence, for the remainder of this paper, we compute utility as the negative perplexity:
@@ 3.2 Computing $\lambda$-similar Labels
We describe our algorithm for identifying $\lambda$-similar labels and their contexts. As we observed before, it is not necessary to iterate over all subsets of the context: it suffices to iterate over all labels below the full context’s label and consider their corresponding subcontexts. Technically, we iterate over the powerset of $\mathcal{P}(\{\bigsqcup_{c\in C^{\prime}}\ell(c)\mid C^{\prime}\subseteq C\})$ of possible output labels and we compute the similarity of the corresponding subcontexts to the full context. Note that the set of labels considered is based on documents in the context; it is finite even when the full lattice of labels $\mathcal{L}$ is infinite (e.g. timestamps).
@@ 1 Function minimal_labels($C,L,x,y$):
2       $\Lambda\leftarrow\emptyset$ for $L^{\prime}\in\textnormal{{children(}}\textnormal{\emph{L}}\textnormal{{)}}$ do
@@ 3.3 Correctness
As utility functions are not necessarily monotonous [39], Algorithm 1 is a heuristic in practice. In Section 5, we evaluate how closely it matches the statement in Proposition 1. Even finding sub-optimal labels can improve over the baseline label propagator.
@@ 3.4 A System for Label Propagation
To integrate Algorithm 1 into existing model architectures and systems, we consider two architectures for augmenting language models with retrieved information: prompt-based retrieval augmentation [22, 16] and $k$NN language models [20].
Prompt-based Augmentation
Augmenting LLM prompts with retrieved documents has become a popular approach to incorporate a non-parametric datastore into an LLM-based pipeline [22, 16]. For example, in a Retrieval Augmented Generation (RAG) setup [22], the documents most relevant for a query are retrieved and added to the context in the prompt to an autoregressive language model.
$k$NN Language Models
In the $k$NN-LM architecture [20], the language model produces an output distribution $p_{\text{LM}}(y,x)$ without taking into account the retrieved documents. A separate distribution $p_{k\text{NN}}(y|x,C)$ is computed based on the retrieved documents only, using $x$ as the criteria for document selection. Finally, a mixture of both distributions (parameterized by a hyperparameter $\gamma$) produces the final retrieval augmented output which is given by
@@
Label Propagation Wrapper
Our approach can be integrated into both of the above architectures as a wrapper around the existing system. We assume that every document $c$ that can be retrieved has a label $\ell(c)$. In practice, if this is not the case, unlabeled documents can be assigned the most restrictive label (i.e. the top of the lattice). The process then proceeds as follows:
@@ 3.5 Computational Cost
The label propagator wrapper based on prompt-based augmentation requires a number of LLM calls that in the worst-case is exponential in the number of documents in the context. This occurs e.g., in a flat bounded lattice where all labels between $\bot$ and $\top$ are incomparable and when each subcontext has a different label.
However, the number of LLM queries is always bounded by the size of the lattice. When the lattice forms a totally ordered set (e.g. two-element lattices distinguishing between trusted vs untrusted or confidential vs public data), Algorithm 1 stops as soon as it finds a subcontext whose utility drops below a $\lambda$ difference w.r.t. the utility of the full context. For richer lattices describing more fine-grained security policies, Algorithm 1 visits a small subset of the lattice in typical queries, either because only a few labels are represented in the context or because the utility drops below the tolerance $\lambda$. Therefore, computational costs are not a major concern for several fundamental lattices. We discuss several optimizations in Section 6.3 that reduce the cost of running the system.
To avoid worst-case costs, alternative algorithms or model architectures may be more suitable. For example, the $k$NN-based architecture we consider in this paper only requires one additional LLM call per query ($p_{\text{LM}}(y|x)$, in addition to the original query $p_{\text{LM}}(y|x,C)$).
@@ 4 Evaluation Setup
Figure 4: Sample documents and QA pairs from the synthetic key-value dataset. The question refers to the social security numbers and dates of birth of person 1 and 2. This information can be obtained by accessing documents $A$, $B$, and $C$ or $A$, and $D$, hence the resulting label of $\{ABC,AD\}$.
@@
[…]
@@ 4.2 Evaluation Metrics
Exact match
The exact match metric computes the average number of completely correct predictions over all questions. That is, for each question we count $1$ if $\Lambda=\Lambda^{\star}$ and 00 otherwise. Exact match gives a direct answer to RQ1, but is very sensitive in that it equally penalizes any imprecision in the label propagator.
Precision and Recall
@@
Model output alignment
The system for label propagation described in Section 3.4 regenerates the output based on the inferred label. While the definition of $\lambda$-similarity guarantees that under the subcontext of the inferred label the original output is almost as likely as in the full context, this does not necessarily guarantee that the regenerated output has good utility w.r.t. to other metrics. To quantify potential deviations, we measure the alignment of the different model outputs w.r.t. the target specified in the dataset i.e., $y^{\star}$ using ROUGE-L [24] metric, which is based on the Longest Common Subsequence (LCS) [17] computation and has been commonly used to estimate translation quality in the past.
In particular, we report the average alignment of the full context output ROUGE-L$(y,y^{\star})$, and the average alignment of the reduced context output after label propagation ROUGE-L$(y^{\prime},y^{\star})$. Furthermore, we report the average difference of the alignment between the reduced context output and full context output i.e., ROUGE-L$(y^{\prime},y^{\star})$ ROUGE-L$(y,y^{\star})$ in two different cases i.e., (i) when label improvement is possible, and (ii) when label improvement is not possible. The difference in output alignment should be nearly zero when label improvement is possible, and highly negative when such improvement is not possible.
@@ 4.3 Baseline
As a baseline for comparison, we use introspection in which the LLM itself is asked to determine which documents in the context were influential. This is inspired by the recent successful application of language models to generate relevant citations for their own outputs [40, 13]. Note that the performance of the introspection technique depends significantly on the prompting technique used, and hence, should be considered as a lower-bound [42].
@@ 4.4 Models
We focus on the Llama-2 model family [41] including its 7B and 70B variants for evaluation. We use instruction-tuned versions of Llama-2 to ensure accurate response generation unless mentioned otherwise. For the $k$NN-LM implementation, we follow [20] and use the model’s penultimate layer representation of the last token (conditioned on all preceding tokens in the document) as the context representation for $k$NN search. Note that while we rely on Llama-2 model family [41] for our experiments, our approach is also applicable to other open-source models, or even proprietary models accessible only via API (though client-side optimizations will be harder to implement).
@@ 4.5 Datasets
Synthetic key-value dataset
@@
News article dataset
Starting from an existing fake news dataset [33], we create pairs of high and low-integrity news articles that discuss similar topics to each other. Contrary to the previous dataset, we focus on a simpler lattice but more complex language in this case. Using GPT-4 [32], we generated QA pairs based on these articles, where some of the answers depended only on the LoInt document, others only on the HiInt document, and some on both documents. An example document pair and QA pair is shown in Figure 5. The dataset in total contains 240 document pairs of corresponding LoInt and HiInt documents, as well as 3465 QA pairs. Compared to a naïve label propagator, it is possible to improve the label on 647 of these QA pairs. Since there is a total order in the lattice, we report label improvement metrics for this dataset.
LLM agent dataset
@@
Motivated by the recent emergence of LLM agents [49] and tool use in LLMs [36], we design an LLM agent dataset that focuses on tool use. Retrieval tools such as web or email search typically return multiple results of varying integrity levels. For example, in the case of web search, official documentation like docs.python.org might be deemed to be higher integrity than community curated sources such as stackoverflow.com. Similarly, emails from verified senders may be deemed to be higher integrity than those from unknown senders.
@@
To cover a wider range of cases, we manually create 40 distinct chat conversations $x$ that involve tool calls. Some of these tools return a set of two documents $C$. We also generate a ground truth model output $y^{\star}$ and label $L^{\star}$. In 20 of the cases, the final label can be improved i.e. $L^{\star}\sqsubset\ell(C)$ due to redundant or irrelevant information within the retrieved LoInt documents. In the remaining 20 cases, the LoInt source is required to produce the output and thus the output label cannot be improved i.e. $L^{\star}=\ell(C)$. We additionally include two in-context examples that are specifically used to specify the output format for the model. Since there is a total order in the lattice, we report label improvement metrics.

### v3

@@ Abstract
Large Language Models (LLMs) are rapidly becoming commodity components of larger software systems. This poses natural security and privacy problems: poisoned data retrieved from one component can change the model’s behavior and compromise the entire system, including coercing the model to spread confidential data to untrusted components. Assuming each piece of information comes with an additional meta-label (such as low/high integrity labels), one promising approach is to tackle this problem at the system level via dynamic information flow (aka taint) tracking. Unfortunately, this approach of propagating the most restrictive input label to the output is too conservative for applications where LLMs operate on inputs retrieved from diverse sources.
In this paper, we propose a novel, more permissive approach to propagate information flow labels through LLM queries. The key idea behind our approach is to propagate only the input labels that were influential in generating the model output and to eliminate the labels of unnecessary inputs. We implement and investigate the effectiveness of two variations of this approach, based on (i) prompt-based retrieval augmentation, and (ii) a $k$-nearest-neighbors language model. We compare these with a baseline that uses introspection to predict the output label. Our experimental results in an LLM agent setting show that our label propagator assigns a more permissive label over the baseline in more than 85% of the cases, which underscores the practicality of our approach.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), Llama (Touvron et al., 2023; Dubey et al., 2024), Mistral (Jiang et al., 2023), and PaLM (Anil et al., 2023) are rapidly becoming commodity components of larger software systems. The inputs to these LLMs often consist of data retrieved from a variety of sources, including websites, productivity software, or tools (Chase, 2022), and their output is usually passed on to other software components for further processing (Xi et al., 2023).
This poses natural security and privacy problems: low integrity inputs (e.g., poisoned data) can change the model’s behavior in unexpected ways and potentially affect the entire system (Greshake et al., 2023; Verma et al., 2024). Similarly, high confidentiality inputs (e.g., confidential documents) can be inadvertently leaked to an untrusted downstream component (Mireshghallah et al., 2023).
One possible approach to address this problem is to rely on the LLM itself for mitigation, for example via introspection of the retrieved inputs or guardrails given in the meta-prompt. However, such defenses can be circumvented with more advanced attacks (Zou et al., 2023; Yuan et al., 2023; Li et al., 2023; Shen et al., 2023), leading to the undesirable cat-and-mouse game that is common in system security. A promising alternative is to tackle this problem at the system level via dynamic information-flow (aka taint) tracking. Information-flow tracking is a standard technique for enforcing integrity and confidentiality properties (Clause et al., 2007; Buiras et al., 2014; Sabelfeld & Myers, 2003), and has been used successfully in many applications, including detecting cross-site-scripting vulnerabilities (Vogt et al., 2007), privacy leaks in mobile applications (Enck et al., 2014) and recently to LLM-based systems (Wutschitz et al., 2023; Wu et al., 2024; Balunovic et al., 2024).
In information-flow tracking, each piece of data is augmented with labels describing its integrity or confidentiality. Labels are usually propagated conservatively: the output of an operation on data is labeled as the most restrictive (i.e., most confidential or least trusted) label of its inputs (Sabelfeld & Myers, 2003). For example, the output of a function that takes two arguments, one trusted and one untrusted, would be labeled as untrusted. A challenge of applying such label propagation mechanisms to LLMs is that the output label would be the upper bound of all inputs (i.e., the context) used for inference. With LLMs having the ability to retrieve documents from different sources, this can quickly become unnecessarily restrictive, a phenomenon known in the literature as label creep (Sabelfeld & Myers, 2003).
@@
To ground our work in a relevant LLM system, we implement and evaluate two different realizations of our label propagator: (i) a prompt-based system (Lewis et al., 2020) where the retrieved documents are provided within the prompt to an autoregressive LLM, and (ii) a kkNN-LM architecture (Khandelwal et al., 2020), where the output distribution is computed as a mixture of the distribution of the model and the retrieved documents.
@@ 2 Problem Setting and Goal
This formulation is sufficiently general to represent many real-world applications of LLMs. For example, in retrieval-augmented generation (RAG) (Lewis et al., 2020; Guu et al., 2020), the context contains documents retrieved from the knowledge-base. Alternatively, if an LLM uses plugins to retrieve external data (e.g., web search, email retrieval, calendar query, etc.) (Schick et al., 2023), the data returned by the plugins is part of the context.
@@ 2.1 Information-Flow Labels
Confidentiality. A canonical example of a security lattice is the set $\{\textnormal{Secret},\textnormal{General}\}$, denoting high and low confidentiality data, where $\textnormal{General}\sqsubseteq\textnormal{Secret}$. The join of Secret and General is Secret (i.e., high confidentiality), but General is the more permissive label.
Integrity. Similarly for integrity, the set $\{\textnormal{HiInt},\textnormal{LoInt}\}$ denotes high and low-integrity data such that $\textnormal{HiInt}\sqsubseteq\textnormal{LoInt}$. The join of LoInt and HiInt is LoInt (i.e., low integrity), but HiInt is the more permissive label.
Product lattices. Figure 2 shows an example of a product lattice for the case of two dimensions: reliability ($\textnormal{HiInt},\textnormal{LoInt}$) and timestamps ($\textnormal{LastMonth},\textnormal{LastWeek},\textnormal{Today}$). The top of the lattice $(\textnormal{LoInt},\textnormal{LastMonth})$ represents the least reliable and least recent documents, while the bottom of the lattice $(\textnormal{HiInt},\textnormal{Today})$ represents the most reliable and most recent documents.
$(\textnormal{LoInt},\textnormal{LastMonth})$$(\textnormal{LoInt},\textnormal{LastWeek})$$(\textnormal{HiInt},\textnormal{LastMonth})$$(\textnormal{LoInt},\textnormal{Today})$$(\textnormal{HiInt},\textnormal{LastWeek})$$(\textnormal{HiInt},\textnormal{Today})$
@@ 3.1 $\lambda$-similar Labels
Our goal is to find labels $L^{\prime}\sqsubseteq L$ such that the output of the language model changes only negligibly when substituting $C$ with $C_{|L^{\prime}}$. We capture “negligible change” by introducing a hyperparameter $\lambda$ and require that the utility of the model’s output under the full context drops by at most $\lambda$ when restricting to the subcontext. Formally:
@@
Definition 1 leaves the choice of the utility function $U$ and the language model $p_{\text{LM}}(y|x,C)$ unspecified, as different applications require custom choices of these functions. In this paper, we focus on language modeling where perplexity is a common way to measure utility (Kaplan et al., 2020). Hence, for the remainder of this paper, we compute utility as the negative perplexity:
@@ 3.2 Computing $\lambda$-similar Labels
We describe our algorithm for identifying $\lambda$-similar labels and their contexts. As we observed before, it is not necessary to iterate over all subsets of the context: it suffices to iterate over all labels below the full context’s label and consider their corresponding subcontexts. Technically, we iterate over the powerset of $\mathcal{P}(\{\bigsqcup_{c\in C^{\prime}}\ell(c)\mid C^{\prime}\subseteq C\})$ of possible output labels and we compute the similarity of the corresponding subcontexts to the full context. Note that the set of labels considered is based on documents in the context; it is finite even when the full lattice of labels $\mathcal{L}$ is infinite (e.g., timestamps).
@@ 1 Function minimal_labels($C,L,x,y$):
2 $\Lambda\leftarrow\emptyset$
3 for $L^{\prime}\in\textnormal{{children(}}\textnormal{\emph{L}}\textnormal{{)}}$ do
@@ 3.3 Correctness
As utility functions are not necessarily monotonous (Shi et al., 2023), Algorithm 1 is a heuristic in practice. In Section 5, we evaluate how closely it matches the statement in Proposition 1. Even finding sub-optimal labels can improve over the baseline label propagator.
@@ 3.4 A System for Label Propagation
To integrate Algorithm 1 into existing model architectures and systems, we consider two architectures for augmenting language models with retrieved information: prompt-based retrieval augmentation (Lewis et al., 2020; Guu et al., 2020) and kkNN language models (Khandelwal et al., 2020).
Prompt-based Augmentation.
Augmenting LLM prompts with retrieved documents has become a popular approach to incorporate a non-parametric datastore into an LLM-based pipeline (Lewis et al., 2020; Guu et al., 2020). For example, in a Retrieval Augmented Generation (RAG) setup (Lewis et al., 2020), the documents most relevant for a query are retrieved and added to the context in the prompt to an autoregressive language model.
kkNN Language Models.
In the kkNN-LM architecture (Khandelwal et al., 2020), the language model produces an output distribution $p_{\text{LM}}(y,x)$ without taking into account the retrieved documents. A separate distribution $p_{k\text{NN}}(y|x,C)$ is computed based on the retrieved documents only, using $x$ as the criteria for document selection. Finally, a mixture of both distributions (parameterized by a hyperparameter $\gamma$) produces the final retrieval augmented output which is given by
@@
Label Propagation Wrapper.
Our approach can be integrated into both of the above architectures as a wrapper around the existing system. We assume that every document $c$ that can be retrieved has a label $\ell(c)$. In practice, if this is not the case, unlabeled documents can be assigned the most restrictive label (i.e., the top of the lattice). The process then proceeds as follows:
@@ 3.5 Computational Cost
The label propagator wrapper based on prompt-based augmentation requires a number of LLM calls that, in the worst case, is exponential in the number of documents in the context. This occurs e.g., in a flat bounded lattice where all labels between $\bot$ and $\top$ are incomparable and when each subcontext has a different label.
However, the number of LLM queries is always bounded by the size of the lattice. When the lattice forms a totally ordered set (e.g. two-element lattices distinguishing between trusted vs. untrusted or confidential vs. public data), Algorithm 1 stops as soon as it finds a subcontext whose utility drops below a $\lambda$ difference w.r.t. the utility of the full context. For richer lattices describing more fine-grained security policies, Algorithm 1 visits a small subset of the lattice in typical queries, either because only a few labels are represented in the context or because the utility drops below the tolerance $\lambda$. Therefore, computational costs are not a major concern for several fundamental lattices. We discuss several optimizations in Section 6.3 that reduce the cost of running the system. Liu et al. (2024a) also proposed tricks to further accelerate the computation of these context attributions.
To avoid worst-case costs, alternative algorithms or model architectures may be more suitable. For example, the kkNN-based architecture we consider in this paper only requires one additional LLM call per query ($p_{\text{LM}}(y|x)$, in addition to the original query $p_{\text{LM}}(y|x,C)$).
@@ 4 Evaluation Setup
Figure 4: Sample documents and QA pairs from the synthetic key-value dataset. The question refers to the social security numbers and dates of birth of person 1 and 2. This information can be obtained by accessing documents $A$, $B$, and $C$ or AA, and $D$, hence the resulting label of $\{ABC,AD\}$.
@@
Sample QA pair from the news article datasetÄpgjy [Question] What are the new fruit-inspired colors for the iPhone lineup mentioned in the article about Apple cutting prices on lower-end iPads and adding colors to the iPhone lineup? [Answer] The new fruit-inspired colors for the iPhone lineup mentioned in the article are Cherry (red), Lemon (yellow), Lime (green), Blueberry (blue), and Grape (purple). [Label] $\{\textnormal{LoInt}\}$ [Question] What is the new color introduced for the iPhone according to the article about Apple cutting prices on lower-end iPads and releasing red iPhones? [Answer] According to the article, the new color introduced for the iPhone is red. [Label] $\{\textnormal{HiInt}\}$
@@ 4.2 Evaluation Metrics
Exact match.
The exact match metric computes the average number of completely correct predictions over all questions. That is, for each question we count $1$ if $\Lambda=\Lambda^{\star}$ and 0 otherwise. Exact match gives a direct answer to RQ1, but is very sensitive in that it equally penalizes any imprecision in the label propagator.
Precision and Recall.
@@
Model output alignment.
The system for label propagation described in Section 3.4 regenerates the output based on the inferred label. While the definition of $\lambda$-similarity guarantees that under the subcontext of the inferred label the original output is almost as likely as in the full context, this does not necessarily guarantee that the regenerated output has good utility w.r.t. to other metrics. To quantify potential deviations, we measure the alignment of the different model outputs w.r.t. the target specified in the dataset i.e., $y^{\star}$ using ROUGE-L (Lin, 2004) metric, which is based on the Longest Common Subsequence (LCS) (Hunt & Szymanski, 1977) computation and has been commonly used to estimate translation quality in the past.
In particular, we report the average alignment of the full context output ROUGE-L(y,y)($y^{\star}$), and the average alignment of the reduced context output after label propagation ROUGE-L(y,y)($^{\prime},y^{\star}$). Furthermore, we report the average difference of the alignment between the reduced context output and full context output i.e., $(y^{\prime},y^{\star}) - ROUGE-L$$(y,y^{\star})$ in two different cases i.e., (i) when label improvement is possible, and (ii) when label improvement is not possible. The difference in output alignment should be nearly zero when label improvement is possible, and highly negative when such improvement is not possible.
@@ 4.3 Baseline
As a baseline for comparison, we use introspection in which the LLM itself is asked to determine which documents in the context were influential. This is inspired by the recent successful application of language models to generate relevant citations for their own outputs (Taylor et al., 2022; Gao et al., 2023). Note that the performance of the introspection technique depends significantly on the prompting technique used, and hence, should be considered as a lower-bound (Turpin et al., 2024).
@@ 4.4 Models
We focus on the Llama-2 model family (Touvron et al., 2023) including its 7B and 70B variants for evaluation. We use instruction-tuned versions of Llama-2 to ensure accurate response generation unless mentioned otherwise. For the kkNN-LM implementation, we follow Khandelwal et al. (2020) and use the model’s penultimate layer representation of the last token (conditioned on all preceding tokens in the document) as the context representation for kkNN search. Note that while we rely on Llama-2 model family (Touvron et al., 2023) for our experiments, our approach is also applicable to other open-source models, or even proprietary models accessible only via API (though client-side optimizations will be harder to implement).
@@ 4.5 Datasets
Synthetic key-value dataset.
@@
News article dataset.
Starting from an existing fake news dataset (Pérez-Rosas et al., 2018), we create pairs of high and low-integrity news articles that discuss similar topics to each other. Contrary to the previous dataset, we focus on a simpler lattice but more complex language in this case. Using GPT-4 (OpenAI, 2023), we generated QA pairs based on these articles, where some of the answers depended only on the LoInt document, others only on the HiInt document, and some on both documents. An example document pair and QA pair is shown in Figure 5. The dataset in total contains 240 document pairs of corresponding LoInt and HiInt documents, as well as 3465 QA pairs. Compared to a naïve label propagator, it is possible to improve the label on 647 of these QA pairs. Since there is a total order in the lattice, we report label improvement metrics for this dataset.
LLM agent dataset.
@@
Motivated by the recent emergence of LLM agents (Xi et al., 2023) and tool use in LLMs (Schick et al., 2023), we design an LLM agent dataset that focuses on tool use. Retrieval tools such as web or email search typically return multiple results of varying integrity levels. For example, in the case of web search, official documentation like docs.python.org might be deemed to be higher integrity than community curated sources such as stackoverflow.com. Similarly, emails from verified senders may be deemed to be higher integrity than those from unknown senders.
@@
To cover a wider range of cases, we manually create 40 distinct chat conversations $x$ that involve tool calls. Some of these tools return a set of two documents $C$. We also generate a ground truth model output $y^{\star}$ and label $L^{\star}$. In 20 of the cases, the final label can be improved, i.e., $L^{\star}\sqsubset\ell(C)$ due to redundant or irrelevant information within the retrieved LoInt documents. In the remaining 20 cases, the LoInt source is required to produce the output and thus the output label cannot be improved, i.e., $L^{\star}=\ell(C)$. We additionally include two in-context examples that are specifically used to specify the output format for the model. Since there is a total order in the lattice, we report label improvement metrics.